# Mamba-ND: Selective State Space Modeling for Multi-Dimensional Data

### 08 Feb 2024 v1

@@ 4.3 Arranging Mamba Layers
In addition to making direct changes to the internal structure of Mamba, one can also change the way in which the layers are organized to achieve multi-directionality. We illustrate these variations in Figure 6.
Alternating-Directional: H+H-W+W-T+T-
keeps the sequential ordering of Mamba layers and changes the direction of SSM in each layer in an alternating fashion. The ordering is H+H-W+W-T+T-
Bi-Directional: [H+H-][W+W-][T+T-]
adopts a design on the block level. In each block, the input is passed to two Mamba layers at opposite directions. The ordering is [H+H-][W+W-][T+T-], where each [.][.][ . ] denotes a bidirectional block consisting of two layers. To avoid confusion, we will explicitly refer to this method as [H+H-][W+W-][T+T-]. The term Bi-Directional will mostly be used for the Bi-SSM layer mentioned in Section 4.2.
Quad-Directional: [H+H-W+W-][T+T-]
builds on top of the [H+H-][W+W-][T+T-] design. It further groups the H and W directions. This design is inspired by works in Video Recognition such as [42], which factorize 3D convolution into a 2D operation on the spatial dimensions and a 1D operation in the temporal domain.
@@ 4.4 Scan Factorization
In order to mitigate the quadratic complexity of the Transformer, prior works [19] factorize full 3D attention into three 1D attentions along each axis. While SSMs already achieve linear complexity, the sequence length is still quadratic in the length of a single dimension. In this work, we also explore various ways of factorizing an SSM scan into multiple smaller scans. For an input array of dimensions $T\times H\times W$, the standard approach is to flatten it into a single sequence of length $THW$. Alternatively, we can factorize it into $T$ sequences with length $HW$, or $TH$ sequences with length $W$. We visualize these factorization techniques in Figure 6. Since SSM only retains one copy of a state per sequence in the linear scan process, increasing the number of sequences in parallel actually leads to an increase in memory consumption and training time because more hidden states need to be materialized in the GPU memory. However, we note that these sub-sequences need not be computed in parallel. Hence a better implementation in the future may reduce the memory cost of this design. Thus, we still investigate the performance of this design.
@@ 4.5 Final Design
After extensive experiments, our final design uses the standard 1D-SSM layer and an alternating-directional: (H+H-W+W-T+T-) at the block level. We find that this simple design surprisingly outperforms more complicated ones. We provide more details of our experiments in Section 5.
Our overall image is shown in Figure 2. Given multi-dimensional input data, we first patchify it into a 1D sequence. During this process, we keep track of the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1-D SSM layers that process the sequence in alternating orderings. In 2D space, we use the order H+H-W+W-. In 3D space, we use the order H+H-W+W-T+T-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream tasks.
@@ 5 Experiments
Settings
We aim to evaluate the effectiveness of Mamba-ND  on various multi-dimensional data tasks. Specifically, we use ImageNet-1K [9] for image classification, HMDB-51 [22] for action recognition, and ERA5 5.625-degree weather forecasting [18] for 3D tasks. ImageNet-1K is a large-scale dataset containing 1.2 million images across 1000 classes, HMDB-51 is an action recognition dataset comprising 7,030 video clips categorized into 51 action classes, and ERA5 consists of 3D atmospheric weather measurements, such as temperature and wind speed, across 13 pressure levels. We use the standard train and validation split for ImageNet, split1 of HMDB-51, and data from the years 1979-2016 for ERA5 as the training set, data from 2017 as the validation set, and data from 2018 as the test set.
@@
We measure top-1 accuracy for image classification and action recognition tasks. For climate forecasting, we report both the Residual Mean Squared Error (RMSE) and the Anomaly Correlation Coefficient (ACC), which is calculated as:
$ACC=\frac{\sum(f-c)(a-c)}{\sqrt{\sum(f-c)^{2}\sum(a-c)^{2}}}$ (7)
Here, $f$ represents the forecast, $c$ is the climatology (long time average measurement), and $a$ is the prediction target.
@@ 5.1 Image Classification
Following the approach of previous studies [12, 33, 17], Mamba-ND  is trained on the ImageNet-1K dataset for 300 epochs using the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $31e-$ 3. The results are presented in Table 1. Our model demonstrates superior performance compared to transformer-based models when operating under similar conditions, and it achieves results on par with the state-of-the-art state-space model, S4ND [12]. We compare our results with Hyena [37], ViT [12], and S4ND [33]. Notably, Mamba-ND-S shows a remarkable improvement of $+1.5$ in accuracy when compared to ViT, while simultaneously reducing the parameter count by 20.7%. This performance gap is consistent with prior research on 1D sequences [15], where Mamba consistently outperforms transformers with fewer parameters.
Table 1: ImageNet 1K Classification Results. We report Top 1 Accuracy on the validation set. Mamba-ND-S shows a remarkable improvement of $+1.5$ in accuracy when compared to ViT-B while reducing the parameter count by 20.7%
@@
Mamba-2D-T 224 24M 79.2
Mamba-2D-S 224 63M 79.4
Mamba-2D-B 224 92M 80.2
@@ 5.2 Video Action Recognition
We sample 32 frames at a frame interval of 2 from each video clip, which amounts to around 2 seconds of video. We use the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $31e-$ 3. The learning rate on the backbone is multiplied by a factor of 0.1. We train our model with a global batch size of 64 for 50 epochs. We select Inception-I3D [4] and ConvNeXt-I3D [27] as our convolutional baseline. For transformer-based models, we select Video Swin Transformer [28].
We report the results in Table 2. Our method achieves a $+0.9$ accuracy compared to Video Swin Transformer while using 39% fewer parameters. Compared to the state-of-the-art ConvNet ConvNeXt, we achieve a $+5.5$ accuracy.
[…]
Model Acc.$\uparrow$ Params Memory Samples/s
ConvNeXt-I3D* 58.1 29M - -
S4ND-ConvNeXt-3D* 62.1 29M - -
Inception-I3D 49.8 25M - -
[…]
Video-Swin-T 53.0 30M 35GB 22.6
Video-Swin-S 58.1 54M 73GB 39.2
Mamba-2D 51.2 24M 12GB 39.2
Mamba-3D 59.0 36M 17GB 19.6
5.3 Global Weather Forecast
For weather forecasting tasks, we use Cli-ViT [34, 35, 36] and Pangu-Weather [3] as our baselines. Because these models were originally trained on terabytes of high-resolution data, typically around 1 degree, for a long time, we cannot directly compare them due to computational constraints. Instead, we use the 5.625-degree version of the ERA data across 7 pressure levels. Our baselines Cli-ViT are based on Vision Transformer, and Pangu-Weather is based on Swin-Transformer. For Cli-ViT, we use the official implementation from ClimaX [34]. For Pangu-Weather, we re-implement it in PyTorch. We train all models from scratch for 50 epochs, using a learning rate of 0.0005. The learning rate follows a cosine decay schedule with a warm-up period of 5 epochs. For Mamba-3D, we use a patch size of $2\times 2\times 2$.
We show the results in Table 3. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
Table 3: ERA5 5.625 Degree Weather Forecast Results. We report the Anomaly Correlation Coefficient (ACC) and RMSE on geopotential at the 500 hPa pressure level. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
@@ 5.3 Global Weather Forecast
Mamba-3D SSM 50M 433 90.1
@@ 5.4 Meta Architectures
We perform extensive ablation studies on various design choices mentioned in Section 4. We show that the alternating-directional design is the simplest and most effective one among a wide range of possible choices.
@@
We use ImageNet-1K and HMDB-51 classification as our benchmarks. We perform an ablation study on Mamba-2D and Mamba-3D. We adopted various designs mentioned in Section 4.2. Results show that our final alternating-directional design achieves stronger performance than all proposed layer-wise changes. We show results in Table 4. Compared with the 1D-Mamba baseline, in which no special design is adopted after naively flattening the multi-dimensional sequence into a 1D sequence, we achieved a $+4.8$ accuracy on ImageNet and a $+26.9$ accuracy on HMDB-51.
Additionally, we observe that most of the proposed multi-directional designs are able to outperform the naive and bi-directional baselines by considerable margins, suggesting the importance of multi-directional design.
@@
We perform comprehensive ablation studies on various organizations of SSM layers mentioned in Section 4.3 at the block level. In addition to three choices discussed in Section 4.3, we also experiment with a hex-directional design [H+H-W+W-T+T-], in which inputs of a block are processed by multiple layers at different ordering in parallel, and the results are summed together in the end. We show the results in Table 5. Contrary to intuition and the layer-level results, adding multi-directional design at the block level degrades the performance. We hypothesize that this behavior is caused by the reduced number of model depths. Since we keep the number of layers fixed across all designs, incorporating some level of multi-directional parallel processing will inevitably reduce the depths of the computation graph. We provide more discussions in the discussion section.
Table 5: Ablation Study on Layer Arrangement. We report top-1 accuracy on the HMDB-51 dataset. H+H-W+W-T+T- is the top-performing design.
Ordering HMDB-51$\uparrow$
@@
5.5 Scan Factorization
While factorizing the sequence in the current implementation of Mamba leads to considerable memory and runtime overhead, these costs are external to the designs themselves and may be patched in the future. For example, a scan of N sequences of length $L$ can be considered as a long scan of one sequence of $NL$ with $\bar{A}$ from Equation 5 set to zero in places where the scan operation moves from one short sequence to another. Hence, we find it useful to study the effects of various factorization techniques despite their inferior runtime efficiency.
We show such results in Table 6. We also report the memory cost on a single Nvidia A100 GPU. While all factorizations lead to worse performance, we find that 2D+1D factorization outperforms the 2D+3D setup, suggesting there may be merits of having certain layers dedicated to processing temporal correspondence. Because the cost of 1D factorization is high, we choose not to explore it further at this time.
Table 6: Ablation Study on Various Factorization Policies + indicates layer ordering. For example, the notation 2D+3D represents a 2D factorization layer and a 3D layer with no factorization arranged sequentially. In all studies, the total number of layers is fixed. We also show the training memory cost and the number of sequences. $B$ is the batch size, and $D$ is the length of a single dimension.
Factorization HMDB-51$\uparrow$ Mem Num.
@@ 5.5 Scan Factorization
5.6 Video Specific Ablations

### 12 Mar 2024 v2

@@ 4.3 Arranging Mamba Layers
In addition to making direct changes to the internal structure of Mamba, one can also change the way in which the layers are organized to achieve multi-directionality. We illustrate these variations in Fig. 4(a).
Alternating-Directional: H+H-W+W-T+T- keeps the sequential ordering of Mamba layers and changes the direction of SSM in each layer in an alternating fashion. The ordering is H+H-W+W-T+T-
Bi-Directional: [H+H-][W+W-][T+T-] adopts a design on the block level. In each block, the input is passed to two Mamba layers at opposite directions. The ordering is [H+H-][W+W-][T+T-], where each [.][.][ . ] denotes a bidirectional block consisting of two layers. To avoid confusion, we will explicitly refer to this method as [H+H-][W+W-][T+T-]. The term Bi-Directional will mostly be used for the Bi-SSM layer mentioned in Sec. 4.2.
Quad-Directional: [H+H-W+W-][T+T-] builds on top of the [H+H-][W+W-][T+T-] design. It further groups the H and W directions. This design is inspired by works in video recognition e.g.,[47], which factorize 3D convolution into a 2D operation on the spatial dimensions and a 1D operation in the temporal domain.
@@ 4.4 Scan Factorization
In order to mitigate the quadratic complexity of the Transformer, prior works [20] factorize full 3D attention into three 1D attentions along each axis. While SSMs already achieve linear complexity, the sequence length is still quadratic in the length of a single dimension. In this work, we also explore various ways of factorizing an SSM scan into multiple smaller scans. For an input array of dimensions $T\times H\times W$, the standard approach is to flatten it into a single sequence of length $THW$. Alternatively, we can factorize it into $T$ sequences with length $HW$, or $TH$ sequences with length $W$. We visualize these factorization techniques in Fig. 4(b). Since SSM only retains one copy of a state per sequence in the linear scan process, increasing the number of sequences in parallel actually leads to an increase in memory consumption and training time because more hidden states need to be materialized in the GPU memory. However, we note that these sub-sequences need not be computed in parallel. Hence a better implementation in the future may reduce the memory cost of this design. Thus, we still investigate the performance of this design.
@@ 4.5 Final Design
After extensive experiments, our final design uses the standard 1D-SSM layer and an alternating-directional: (H+H-W+W-T+T-) at the block level. We find that this simple design surprisingly outperforms more complicated ones. We provide more details of our experiments in Sec. 5.
Our overall image is shown in Fig. 2. Given multi-dimensional input data, we first patchify it into a 1D sequence. During this process, we keep track of the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1-D SSM layers that process the sequence in alternating orderings. In 2D space, we use the order H+H-W+W-. In 3D space, we use the order H+H-W+W-T+T-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream tasks.
@@ 5 Experiments
Datasets and Setups.
We aim to evaluate the effectiveness of Mamba-ND  on various multi-dimensional data tasks. Specifically, we use ImageNet-1K [8] for image classification, HMDB-51 [23] and UCF-101 [43] for action recognition, ERA5 5.625-degree for weather forecasting [19] and BTCV[24] for 3D segmentation. ImageNet-1K is a large-scale dataset containing 1.2 million images across 1000 classes, HMDB-51 and UCF-101 are action recognition datasets comprising 7,030 and 13,320 video clips respectively, BTCV consists of abdominal CT scans of 30 subjects, among which 6 are selected as validation set. ERA5 consists of 3D atmospheric weather measurements, such as temperature and wind speed, across 13 pressure levels. We use the standard train and validation split for ImageNet, split1 of HMDB-51, and data from the years 1979-2016 for ERA5 as the training set, data from 2017 as the validation set, and data from 2018 as the test set.
@@
We measure top-1 accuracy for image classification and action recognition tasks. For weather forecasting, we report both the Residual Mean Squared Error (RMSE) and the Anomaly Correlation Coefficient (ACC). For 3D CT segmentation, we report the DICE score.
@@ 5.1 Image Classification
Following the approach of previous studies [11, 37, 18], Mamba-ND  is trained on the ImageNet-1K dataset for 300 epochs using the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $31e-$ 3. We use a patch size of 8. The results are presented in Tab. 1. Our model demonstrates superior performance compared to transformer-based models when operating under similar conditions, and it achieves results on par with the state-of-the-art state-space model, S4ND [11]. We compare our results with Hyena [41], ViT [11], and S4ND [37]. Notably, Mamba-ND-S shows a remarkable improvement of $+3.8$ in accuracy when compared to ViT, while simultaneously reducing the parameter count by 20.7%. This performance gap is consistent with prior research on 1D sequences [14], where Mamba consistently outperforms transformers with fewer parameters.
Table 1: ImageNet 1K Classification Results. We report Top 1 Accuracy on the validation set. Mamba-ND-S shows a remarkable improvement of $+3.8$ in accuracy when compared to ViT-B while reducing the parameter count to 20.7%.
@@
DeiT-S 224 22M 79.8
DeiT-B 224 86M 81.8
Swin-T 224 28M 81.3
Swin-B 224 88M 83.5
Mamba-2D-S 224 24M 81.7
Mamba-2D-B 224 92M 83.0
[…]
Model HDMB-51 $\uparrow$ UCF-101 $\uparrow$ Params Memory Samples/s
ConvNeXt-I3D* 58.1 - 29M - -
S4ND-ConvNeXt-3D* 62.1 - 29M - -
Inception-I3D 49.8 84.5 25M - -
[…]
Video-Swin-T 53.0 88.3 30M 35GB 22.6
Video-Swin-S 58.1 88.7 54M 73GB 39.2
Mamba-2D 51.2 84.7 24M 12GB 39.2
Mamba-3D 60.9 89.6 36M 17GB 19.6
@@ 5.2 Video Action Recognition
We sample 32 frames at a frame interval of 2 from each video clip, which amounts to around 2 seconds of video. We use the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $45e-$ 4. The learning rate on the backbone is multiplied by a factor of 0.1. We train our model with a global batch size of 64 for 50 epochs. We select Inception-I3D [4] and ConvNeXt-I3D [31] as our convolutional baseline. For transformer-based models, we select Video Swin Transformer [30]. We report the results in Tab. 2. Our method achieves a $+2.8$ accuracy compared to Video Swin Transformer while using only 39% of the parameters.
5.3 Global Weather Forecasting
For weather forecasting, we use Cli-ViT [38, 39, 40] and Pangu-Weather [3] as our baselines. Because these models were originally trained on terabytes of high-resolution data, typically around 1°$\degree$, for a long time, we cannot directly compare them due to computational constraints. Instead, we use the 5.625°$\degree$ version of the ERA data across 7 pressure levels. Our baselines Cli-ViT are based on Vision Transformer, and Pangu-Weather is based on Swin-Transformer. For Cli-ViT, we use the official implementation from ClimaX [38]. For Pangu-Weather, we re-implement it in PyTorch. We train all models from scratch for 50 epochs, using a learning rate of 0.0005. The learning rate follows a cosine decay schedule with a warm-up period of 5 epochs. For Mamba-3D, we use a patch size of $2\times 2\times 2$. We show the results in Tab. 3. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
Table 3: ERA5 5.625°normal-$\degree$ Weather Forecasting Results. Anomaly Correlation Coefficient (ACC) and RMSE on geopotential at the 500 hPa level. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
@@ 5.3 Global Weather Forecasting
Mamba-3D Mamba 50M 433 90.1
5.4 3D Medical Image Segmentation
We evaluate our model on BTCV[24] dataset. BTCV consists of abdominal CT scans of 30 subjects. Each CT scan consists of 80 to 225 slices with 512×512 pixels. We use the report numbers on split0 provided by UNETR with 6 validation samples and 24 training samples. We compare against ViT-Based method (UNETR) [17] and Swin-Based method (Swin-UNETR) [16]. We keep the decoder of UNETR unchanged and only replaced its ViT backbone. We use a patch size of $16\times 16\times 16$. All models have 12 layers. The Small, Small+, and Base variant of Mamba-3D-UNETR has hidden sizes of 384, 512 and 768 respectively. We train all models on a single GPU for 5000 epochs. Memory cost are reported according. We report the DICE scores. Notably, Mamba-3D-S achieves $+0.1$ gain while reducing the number of parameters by 64% when compared with UNETR.W Mamba-3D-B, which has a similar parameter count to UNETR, achieves a gain of +2.7 compared with UNETR.
Table 4: 3D Segmentation on BTCV Dataset. We report the score. Mamba-3D consistently outperforms baseline architectures at with less parameter count.
[…]
UNETR ViT 101M 81.4 10G [16, 16, 16]
Swin-UNETR Swin 63M 83.3 21G [2, 2, 2]
Mamba-3D-UNETR-S Mamba 36M 81.5 8G [16, 16, 16]
Mamba-3D-UNETR-S+ Mamba 55M 83.6 9G [16, 16, 16]
Mamba-3D-UNETR-B Mamba 107M 84.1 9G [16, 16, 16]
@@ 5.5 Meta Architectures
We perform extensive ablation studies on various design choices mentioned in Sec. 4. We show that the alternating-directional design is the simplest and most effective one among a wide range of possible choices.
@@
We use ImageNet-1K and HMDB-51 classification as our benchmarks. We perform an ablation study on Mamba-2D and Mamba-3D. We adopted various designs mentioned in Sec. 4.2. Results show that our final alternating-directional design achieves stronger performance than all proposed layer-wise changes. We show results in Tab. 5. Compared with the 1D-Mamba baseline, in which no special design is adopted after naively flattening the multi-dimensional sequence into a 1D sequence, we achieved a $+4.8$ accuracy on ImageNet and a $+26.9$ accuracy on HMDB-51. Additionally, we observe that most of the proposed multi-directional designs are able to outperform the naive and bi-directional baselines by considerable margins.
@@
We perform comprehensive ablation studies on various organizations of SSM layers mentioned in Sec. 4.3 at the block level. In addition to three choices discussed in Sec. 4.3, we also experiment with a hex-directional design [H+H-W+W-T+T-], in which inputs of a block are processed by multiple layers at different ordering in parallel, and the results are summed together in the end. We show the results in LABEL:tab:ordering_results. Contrary to intuition and the layer-level results, adding multi-directional design at the block level degrades the performance. We hypothesize that this behavior is caused by the reduced number of model depths. Since we keep the number of layers fixed across all designs, incorporating some level of multi-directional parallel processing will inevitably reduce the depths of the computation graph. We provide more discussion in Sec. 6.
Table 6: Ablation Studies
[…]
@@
[…]
@@
5.6 Scan Factorization
While factorizing the sequence in the current implementation of Mamba leads to considerable memory and runtime overhead, these costs are external to the designs themselves and may be patched in the future. For example, a scan of N sequences of length $L$ can be considered as a long scan of one sequence of $NL$ with $\bar{A}$ from Eq. 5 set to zero in places where the scan operation moves from one short sequence to another. Hence, we find it useful to study the effects of various factorization techniques despite their inferior runtime efficiency.
We show such results in LABEL:tab:factorization_results. We also report the memory cost on a single Nvidia A100 GPU. While all factorizations lead to worse performance, we find that 2D+1D factorization outperforms the 2D+3D setup, suggesting there may be merits of having certain layers dedicated to processing temporal correspondence. Because the cost of 1D factorization is high, we choose not to explore it further at this time.
6 Discussion
6.1 Effective Receptive Field (ERF)
We visualize the effective receptive field (ERF) [33] of Mamba-ND  in Fig. 6. ERF is computed by setting the gradient of the center patch to one and backpropagating through the network. In particular, we visualize the ERF of ImageNet-1K pretrained weights of Mamba-ND  against the Bi-directional and Uni-directional baseline (Tab. 5, row 5 and Tab. 5, row 4). Sensitivity is measured by the whiteness of a pixel in the visualization. The Uni-Directional model exhibits a sharp cutoff and is insensitive to all patches after it in the flattened sequence. This is clearly undesirable in data where causal relations do not exist. The Bi-Directional model exhibits a global receptive field, but it is heavily biased in the horizontal direction. This is also undesirable because images can have vertical structures. The Multi-Directional model exhibits a more uniform pattern in the sensitivity visualization, which explains their superior performance.
6.2 Depths versus Widths
In Sec. 5.5, we showed that a complicated multi-directional design may not necessarily lead to improved performance over the alternating-directional baseline. We hypothesize that this result is related to the effective depth of the model.
7 Conclusion
Transformers have been the go-to choice in image, language, and video tasks in recent years, just like ConvNets and RNNs in earlier years. Mamba [35] presents itself as a competitive challenger to Transformers in 1D sequence modeling. In this work, we proposed Mamba-ND which successfully extends the strong performance of Mamba to multi-dimensional inputs. Mamba-ND outperforms Transformers on a variety of tasks with significantly fewer parameters, while incurring only subquadratic complexity. Through extensive experiments with alternative designs, we demonstrate the importance of our multi-directional design choices over uni-directional and bi-directional baselines.
8 Acknowledgement
We are grateful to Microsoft Research for supporting this research through their Accelerate Foundation Models Research program. We also thank Schmidt Sciences, Google, and Cisco for their support.
Appendix
9 Additional Ablation Studies
9.1 Additional ImageNet and HMDB-51 Results.
HMDB-51 and UCF-101 training requires ImageNet pretrained weights. Due to architectural differences, different HMDB-51 experiments may load different ImageNet pretrained weights. For completeness, we provide full results indicating which weights are loaded for each experiment. All models are trained with a learning rate of 1e-3 and a patch size of 16 for both pretraining and finetuning. We report the results in Tab. 7.
In addition to the results already referred to in the main paper, we also incorporate additional experiments on applying an alternating design to Bi-SSM. In particular, the default Bi-SSM design processes each layer in $L+$ and $L-$ direction, which is equivalent to $H+$ and $H-$. We experiment with processing the input in $H+$ and $H-$ direction in odd layers and in $W+$ and $W-$ direction in even layers. This leads to a considerable improvement in both ImageNet-1K (+1.5) and HMDB-51 (+8.4), highlighting the effectiveness of alternating directional designs.
We also tried to train models from scratch on HMDB-51. These models all converge very slowly or diverge. We failed to obtain useful results.
Table 7: Detailed Ablation Results of various combinations of design choices. In the 2D Layer column, we list the layer-level design for ImageNet pretraining. In the 3D Layer column, we list the layer-level design for finetuning. (.) denotes layer-level grouping. For Bi-SSM, each layer incorporates both the forward direction and reverse direction, hence no +/+/-+ / - distinction is needed.
[…]
1D-SSM L+ 76.4 L+ 34.9
1D-SSM (H+H-)(W+W-) 76.5 (H+H-)(W+W-)(T+T-) 49.8
1D-SSM H+H-W+W- 79.4 (H+H-W+W-)(T+T-) 47.4
H+H-W+W-T+T- 59.0
1D-SSM (L+L-) 76.3 (L+L-) 46.3
Bi-SSM* L 74.6 L 32.1
Bi-SSM* HW 76.1 HWT 40.5
ND-SSM - 77.2 - 46.7
Multi-Head - 77.6 - 51.5
9.2 Video-Specific Ablations
9.3 $\Delta$ Initialization